hi, I am wondering if you can tell me if I go 120km/h how long will it take me to reach 4.25 km
Answer by jim_thompson5910(35256)

D = r*t

4.25 = 120*t

4.25/120 = t

0.03541667 = t

t = 0.03541667

It will take 0.03541667 hours (approximately)

0.03541667*60 = 2.125 minutes

It will take 2.125 minutes

Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim